Assume an individual’s preference for Fanta (Good x) and Fritters (Good y) are given by


the following utility function: U(x,y)=20= x^0.5y^0.5


REQUIRED:


i) Derive the equation for the indifference curve for this utility function


1
Expert's answer
2021-11-08T09:09:40-0500

"x^{0.5}y^{0.5}=20"

"\sqrt{x}\sqrt{y}=20\\\sqrt{y}=\frac{20}{\sqrt{x}}\\y=(\frac{20}{\sqrt{x}})^2=\frac{400}{x}\\y=\frac{400}{x}"